Systems Staff Meeting Notes 08/05/02

Systems Staff Meeting Notes
August 5, 2002

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Harris, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
  1. The back room in Geoffrey’s house flooded. There’s an old septic tank that needs to be filled in and lines for city water established.
  2. Visitor – Bill Hanssen from Facilities is here taking staff photos for the security system. The “Ying Tang Gang” (Ying, her husband & their realtor) was discovered entering Systems after hours.
  3. Dissertation receiving – we had someone trying to turn in a dissertation to Preservation on Friday and no one was in their office so they came over to Systems. Bobbie accepted the dissertation and wrote the person a receipt. Bill spoke with Cathy Mook about it, she said that the whole process will be going online soon and the dissertations will be delivered directly to the grad school. In the meantime, if it happens again, just take the dissertation and assure them it will be delivered to Preservation as soon as possible.
  4. We’re still on for lunch today at noon – and Martha will be in attendance.
  5. Martha and her husband will be bringing in their machine again – they are ready for it to be reformatted.
Review Minutes of 7/29/02
  1. Announcement #4 should be $500,000 (not $500) for money held back.
  2. Nets #3 – indeed, Bobbie has been throwing out tons of old media.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. No major orders in the queue.
  2. Music – barcode readers – received two. Bobbie is concerned that the readers are not lasting long enough. They seem to die immediately after the 2-year warranty is up.
Support Issues and Liaisons
  1. Xerox – visited Will today – they are looking for a starting point – Will directed them to West. They will hook up one release station and one printer with a laptop (server) in West to test.
  2. Screen saver/workstation lockup to be set at 5 minutes by liaisons.
  1. File moving – NT2R – contains user folders, but is a copied version on a new drive. Old NT2R is still there – read-only. Admin is currently on 3 servers – need to work with Misty to move them. John Ashcraft thought NT5 folders had been moved – not yet. Need to free up space on NT2, 3, & 5 soon.
  2. Security Servers – are being installed – need documentation on BellStor HD.
Software Projects Status
  1. Hardware/Software online requests – ready for Bill’s review.
  2. Software Programming online requests – Cynthia is still working on it.
  3. Ariel to SQL server - Access Services & ILL are trying to get together on LAD stuff (currently it’s scanned at LAD – transferred to the Kangaroo drive – brought back to ILL and converted for Ariel). They feel the process is too complicated. Ariel does TIFF files only, it won’t import non-TIFF requests from other institutions.
  4. Hilton is finished with IRDB mods, needs Bobbie or Geoffrey to sign off on it. This version does not include the sort mod.
  5. Ying – DLC – finished their request from the last review. She told them that Stage One is finished and wrote a report for Erich. Now working on the Gift Opportunity database. Next – Sanborn Map database. The search toll is supposed to be done by FCLA – with our help. Bill thinks we’ll end up doing it all.
  6. Resource Services – they are trying to hire an Electronic Technical Service Coordinator – CLARR type programming interface to Cataloging. Winston is chairing the committee.
Web Projects
  1. The small committee has not met yet.

Back to the Systems Homepage

Last updated August 12, 2002
by Debra Harris